OpenStudy (anonymous):

If dy/dx=1/x, then the average rate of change of y with respect to x on the closed interval [1, 4] is a) -1/4 b) (1/2)*ln(2) c) (2/3)*ln(2) d) 2/5 e) 2 Answer: C

OpenStudy (anonymous):

@eliassaab @thomaster @ganeshie8

OpenStudy (anonymous):

@whpalmer4 @AravindG

OpenStudy (anonymous):

@kirbykirby @dumbcow

OpenStudy (anonymous):

if \(\frac{dy}{dx}=\frac{1}{x}\) then \(y=\ln(x)+c\)

OpenStudy (anonymous):

\[\frac{\ln(4)-\ln(1)}{4-1}=\frac{\ln(4)-0}{3}=\frac{\ln(4)}{3}\]

OpenStudy (anonymous):

Thank you everyone!

OpenStudy (anonymous):

\[\frac{dy}{dx}=\frac{1}{x}\] \[dy=\frac{1}{x}dx\] \[\int dy=\int\frac{1}{x}dx\] \[\int\limits_{0}^{y}dy=\int\limits_{1}^{4}\frac{1}{x}dx\] \[y=ln(x)|_1^4\] \[y=ln(4)-ln(1)\] \[y=ln(2^2)-0\] \[\boxed{\boxed{y=2ln(2)}}\]

OpenStudy (anonymous):

Thank you for the help!